Insurance companies are interested in knowing the population percent of drivers who always buckle up before riding in a car. They randomly survey 380 drivers and find that 288 claim to always buckle up. Construct a 97% confidence interval for the population proportion that claim to always buckle up Box 1: Enter your answer using interval notation. Example: [2.1,5.6172) Use U for union to combine intervals. Example: (-o0,2] U [4.00) Enter DNE for an empty set. Use oo to enter Infinity License Points possible: 1 Unlimited attempts.
Please show steps in IT84 calculator please.
Answer #1

A (1-alpha)% for population proportion is given by

p in ( widehat{p} - Z_{alpha /2}*sqrt{widehat{p}*(1-widehat{p})/n} , widehat{p} + Z_{alpha /2}*sqrt{widehat{p}*(1-widehat{p})/n} )

where widehat{p} = X/n , X = number of drivers who claim to buckle up and n = sample size , Q0,03

Hence X= 288 , n = 380 and widehat{p} = 288/380 = 0.758 and Z_{alpha /2} = Z 0.015 = 2.1701.

Hence p in ( 0.758 - 2.1701 * V0.758 * 0.242/380 , 0.758 +2.1701 * V0.758 * 0.242/380 )

or p in ( 0.710 , 0.806 )

We have used simple arithmetic functions and square root function along with values from normal probability tables to find out that a 97% confidence interval for the population proportion that claim to always buckle up is (0.710, 0.806)